# Gene-DML: Dual-Pathway Multi-Level Discrimination for Gene Expression Prediction from Histopathology Images

### v1

@@ 4.1 Dataset and Implementation Details
In this study, we conducted experiments on three ST datasets that are collected in the HEST-1k [14] benchmark: HER2ST [2], STNet [10], and skinST [15]. Specifically, two breast cancer ST datasets – HER2ST, which focuses exclusively on HER2-positive breast tumor subtypes, and STNet, which includes a broader range of breast cancer subtypes – provide approximately $\mathbf{13,620}$ and $\mathbf{68,050}$ spatially resolved spots, respectively. Additionally, skinST contains approximately $\mathbf{23,205}$ spots from spatial profiling of skin tissue sections. Each spot comprises its coordinates on the WSI sample and the corresponding gene expression profiles with around $20,000$ genes for breast and skin cancer. To prevent data leakage, we ensure that samples from the same patient do not appear in both the training and testing datasets. Following [6], we adopt the one-patient-one-fold cross-validation (CV) approach for the HER2ST (8-fold) and skinST (4-fold) datasets; for STNet, which comprises $68$ samples from $23$ patients, we perform an 8-fold CV, ensuring that all samples from the same patient are assigned to the same fold. To evaluate the generalization performance of Gene-DML, we perform external tests on three Visium datasets from 10x Genomics from independent breast cancer patients. In our experiments, we use Pearson Correlation Coefficient (PCC), Mean Squared Error (MSE), and Mean Absolute Error (MAE) as evaluation metrics:
@@
At the data-preprocessing stage, we crop patches with $224\times 224$ pixel patches centered on each spatial spot from WSIs to construct image tiles. For gene expression profiles, we select 250 genes per dataset (details refer to the supplementary), applying spot-wise normalization followed by log transformation according to [10]. For feature grouping, we utilize K-means [36] method. Gene-DML is optimized using the Adam optimizer with an initial learning rate of $1e^{-4}$. The learning rate is decayed by a factor of $0.95$ every $20$ epochs using a StepLR scheduler. The batch size is set to $256$ during training, and we report the mean of the model’s performance evaluated on all spots in every validation WSI during testing. All the experiments are trained on an NVIDIA RTX A6000 (48GB) GPU.
@@ 4.2 Cross-validation Performance of Gene-DML
To evaluate the performance of Gene-DML, we conducted cross-validation experiments on the public ST datasets, HER2ST, STNet, and skinST, and compared the spatial gene expression result predicted by Gene-DML with recent-year existing methods as baselines: ST-Net [10], EGN [34], BLEEP [33], HistoGene [21], Hist2ST [37], TRIPLEX [6], and M2OST [29]. We re-implement the baseline methods by their public code repositories under the same CV fold splitting settings on the same datasets.
Model 10X Visium-1 10X Visium-2 10X Visium-3
@@
ST-Net [10] 20’ 0.423 0.505 -0.026 -0.000 0.395 0.492 0.091 0.193 0.424 0.508 -0.032 0.008
EGN [34] 23’ 0.421 0.512 0.003 0.024 0.328 0.443 0.102 0.157 0.303 0.425 0.106 0.220
BLEEP [33] 23’ 0.367 0.470 0.106 0.221 0.289 0.406 0.104 0.260 0.298 0.415 0.114 0.229
TRIPLEX [6] 24’ 0.351 0.464 0.136 0.241 0.282 0.407 0.155 0.356 0.285 0.410 0.118 0.282
@@
In comparison with existing works, as shown in Tab. 1, Gene-DML consistently achieves the best performance across all metrics, MSE, PCC(A), and PCC(H), demonstrating its superiority in accurately capturing spatial gene expression patterns. Specifically, for HER2ST dataset, Gene-DML obtains the lowest MSE of $0.210$, outperforming the second-best TRIPLEX ($0.228$) by $0.018$. In terms of PCC(A), our method achieves $0.331$, which surpasses TRIPLEX ($0.314$) by $+0.017$, and shows the most significant gain in PCC(H) at $0.541$, improving by $+0.044$ over TRIPLEX ($0.497$). As for STNet dataset, Gene-DML achieves the best MSE of $0.179$, a $0.002$ reduction from Hist2ST ($0.181$), and attains a notable gain in PCC(A) with $0.237$, improving upon TRIPLEX ($0.206$) by $+0.031$. In high-predictive gene correlation (PCC(H)), our method leads $0.384$, ahead $0.353$ of M2OST by $+0.031$. On the skin dataset skinST, Gene-DML achieves the most substantial improvements. Its MSE of $0.237$ is better than $0.268$ of TRIPLEX by $+0.031$. In PCC(A), it scores $0.433$, significantly outperforming the second-best TRIPLEX ($0.374$) by $+0.059$. For PCC(H), Gene-DML leads with $0.520$, an improvement of $+0.030$ over M2OST and TRIPLEX ($0.490$). Due to space limitations, all MAE results are provided in the supplementary material.
@@ 4.3 Generalization Performance of Gene-DML
Tab. 2 presents the generalization performance of Gene-DML compared to recent existing works across three unseen Visium ST datasets. The results are produced by using the pre-trained Gene-DML model on the breast cancer dataset (HER2ST). Our method achieves the highest Pearson Correlation Coefficient (PCC) on both PCC(A) and PCC(H) genes across all datasets, highlighting its superior ability to capture biologically meaningful patterns that generalize beyond training distributions. Specifically, Gene-DML obtains $0.160$, $0.192$, and $0.155$ for metric PCC(A) with an improvement from TRIPLEX by $+0.024$, $+0.037$, and $+0.037$ on data 10X Visium-1, -2, and -3, respectively; notably, Gene-DML also achieves a new state-of-the-art PCC(H) of $0.415$, $0.411$, and $0.393$ on 10X Visium-1, -2, and -3, respectively, surpassing all prior approaches. While our MSE and MAE scores are not always the best, this can be attributed to Gene-DML’s emphasis on semantic alignment and structure-preserving embeddings rather than purely minimizing per-spot regression loss. The slight trade-off in MSE/MAE is acceptable, as PCC offers a more interpretable measure of functional correlation, especially for HPGs critical in downstream biological interpretation.
@@ 4.4 Visualization of Cancer Biomarker Genes
GNAS [1] and FASN [16] are two of the highly predictive genes across all datasets we utilized, and well-established cancer biomarker genes. Fig. 3 presents a qualitative comparison of gene expression prediction with indicating PCC values for GNAS and FASN across two representative tissue WSIs from various patients, with the original WSI, gene expression ground truth, TRIPLEX prediction, and Gene-DML prediction. Specifically, each hexagonal tile corresponds to a spatial transcriptomics (ST) spot, with its color indicating the predicted gene expression level, normalized to the range $[0,1]$. Gene-DML consistently outperforms the model TRIPLEX with the second-best quantitative results in Tab. 1 in accurately recovering the spatial expression patterns, as evidenced by the visibly sharper regional expression and higher PCC values (e.g., $0.608$ vs. $0.440$ and $0.846$ vs. $0.813$ for GNAS; $0.772$ vs. $0.621$ and $0.810$ vs. $0.760$ for FASN) for both relatively hard (sample #1) and easy (sample #2) cases. The robust prediction of GNAS and FASN not only reflects the strength of features learned by the dual multi-scale discrimination but also is widely implicated in tumorigenesis and cancer progression, further underscoring the biological relevance and clinical utility of the Gene-DML framework for spatial biomarker discovery. Visualizations of additional samples are provided in the supplementary material.
@@
Table 3: Ablation study on framework components on the dataset skinST. Abbreviations in this table include instance-level discrimination (I-L Dis.) and instance-group discrimination (I-G Dis.).
@@ 4.5 Ablation Study
We performed extensive experiments to assess the effectiveness of each framework component and values of hyperparameters $k$ (feature group size), $\lambda$ (the weight of instance-group alignment loss), and $\tau_{I-G}$ (the temperature of instance-group similarity calculation) on skinST dataset.
We report the ablation results on each key component of Gene-DML, as illustrated in Tab. 3. It highlights the effectiveness of the proposed multi-scale instance-level and cross-level instance-group discrimination modules. Removing both components, a plain bimodal framework, yields the weakest performance. Incorporating the multi-scale instance-level discrimination alone improves performance $0.015$ on MSE, $+0.039$ on PCC(A), and $+0.053$ on PCC(H). Similarly, adding cross-level instance-group discrimination alone yields a gain $0.017$ on MSE, $0.051$ on PCC(A), and $0.082$ on PCC(H) from the plain bimodal framework. When both modules are integrated, Gene-DML achieves the best performance across all metrics, showing a cumulative improvement of $0.060$ MSE, $+0.164$ PCC(A), and $+0.124$ PCC(H) over the baseline. These results validate the complementary benefits of a dual-pathway multi-level discrimination strategy.
@@
[…]
@@ 5 Conclusion
Appendix B: Additional Ablation Study
@@
Appendix C: Additional Visualization
@@
Appendix D: Additional Implementation Details
@@ Appendix A Additional Cross-validation Performance
Appendix B Additional Ablation Study
To evaluate the dual-pathway multi-level discrimination framework, Gene-DML, we perform an ablation study by adding each discrimination pathway sequentially and report the results of MSE, MAE, PCC(A), and PCC(H) metrics. Tab. 8 and Tab. 12 demonstrate the effectiveness of each component in Gene-DML on datasets HER2ST and STNet. They show that, by removing both pathways, a plain bimodal framework yields the weakest performance. By incorporating the multi-scale instance-level discrimination pathway and cross-level instance-group discrimination pathway, respectively, a significant improvement is achieved.
@@ Appendix B Additional Ablation Study
Table 8: Ablation study on components of the dataset HER2ST. Abbreviations in this table include instance-level discrimination (I-L Dis.) and instance-group discrimination (I-G Dis.).
@@
Table 12: Ablation study on components of the dataset STNet. Abbreviations in this table include instance-level discrimination (I-L Dis.) and instance-group discrimination (I-G Dis.).
@@
Appendix C Additional Visualizations
In this section, we provide additional visualizations of the gene expression prediction results, which are normalized into the range $[0,1]$, of the breast cancer biomarker genes GNAS and FASN. Fig. 4 demonstrates the spatial gene expression prediction for GNAS and FASN on the HER2ST dataset. Each row displays a WSI sample alongside the ground-truth expression map and predictions from TRIPLEX (the second-best result method evaluated in Tab. 1 of Sec. 4) and the proposed Gene-DML. Across both genes and multiple WSI samples, Gene-DML consistently produces spatial expression patterns that are visually closer to the ground truth and quantitatively higher in performance, as reflected by the reported Pearson correlation coefficients in parentheses. The same conclusion holds for Fig. 5, which shows the spatial gene expression prediction results on the dataset STNet, further validating the strong gene expression prediction capacity of Gene-DML.
@@ Appendix C Additional Visualizations
Appendix D Additional Implementation Details
To ensure fairness, following TRIPLEX [6], we utilize the various $250$ selected genes for each dataset, i.e., HER2ST, STNet, and skinST, that are related to morphology in histopathology to evaluate the performance of gene expression prediction. For the results of the MSE, MAE, and PCC(A) metrics in the cross-validation experiments, the same $250$ genes were utilized for both the training and testing processes. Fig. 6, Fig. 7, and Fig. 8 demonstrate the $250$ genes exploited to predict gene expression profiles on datasets HER2ST, STNet, and skinST, respectively.

### v3

@@ 4.1 Dataset and Implementation Details
In this study, we conducted experiments on three ST datasets that are collected in the HEST-1k [Jaume_2025] benchmark: HER2ST [Andersson_2021], STNet [he_2020], and skinST [ji_2020]. Specifically, two breast cancer ST datasets – HER2ST, which focuses exclusively on HER2-positive breast tumor subtypes, and STNet, which includes a broader range of breast cancer subtypes – provide approximately $\mathbf{13,620}$ and $\mathbf{68,050}$ spatially resolved spots, respectively. Additionally, skinST contains approximately $\mathbf{23,205}$ spots from spatial profiling of skin tissue sections. Each spot comprises its coordinates on the WSI sample and the corresponding gene expression profiles with around $20,000$ genes for breast and skin cancer. To prevent data leakage, we ensure that samples from the same patient do not appear in both the training and testing datasets. Following [chung_2024], we adopt the one-patient-one-fold cross-validation (CV) approach for the HER2ST (8-fold) and skinST (4-fold) datasets; for STNet, which comprises $68$ samples from $23$ patients, we perform an 8-fold CV, ensuring that all samples from the same patient are assigned to the same fold. To evaluate the generalization performance of Gene-DML, we perform external tests on $3$ breast Visium data and $2$ skin Visium data. In our experiments, we use Pearson Correlation Coefficient (PCC), Mean Squared Error (MSE), and Mean Absolute Error (MAE) as evaluation metrics:
@@
At the data-preprocessing stage, we crop patches with $224\times 224$ pixel patches centered on each spatial spot from WSIs to construct image tiles. For gene expression profiles, we select 250 genes per dataset (details refer to the supplementary), applying spot-wise normalization followed by log transformation according to [he_2020]. For feature grouping, we utilize K-means [Ye_2007] method. Gene-DML is optimized using the Adam optimizer with an initial learning rate of $1e^{-4}$. The learning rate is decayed by a factor of $0.95$ every $20$ epochs using a StepLR scheduler. The batch size is set to $256$ during training, and we report the mean of the model’s performance evaluated on all spots in every validation WSI during testing. All the experiments are trained on an NVIDIA RTX A6000 (48GB) GPU.
@@ 4.2 Cross-validation Performance of Gene-DML
To evaluate the performance of Gene-DML, we conducted cross-validation experiments on the public ST datasets, HER2ST, STNet, and skinST, and compared the spatial gene expression result predicted by Gene-DML with recent-year existing methods as baselines: ST-Net [he_2020], EGN [Yang_2023], BLEEP [xie_2024], HistoGene [pang_2021], Hist2ST [zeng_2022], TRIPLEX [chung_2024], and M2OST [wang2025m2ost]. We re-implement the baseline methods by their public code repositories under the same CV fold splitting settings on the same datasets.
Model Breast Visium-1 Breast Visium-2 Breast Visium-3
@@
ST-Net [he_2020] 20’ 0.423 0.505 -0.026 -0.000 0.395 0.492 0.091 0.193 0.424 0.508 -0.032 0.008
EGN [Yang_2023] 23’ 0.421 0.512 0.003 0.024 0.328 0.443 0.102 0.157 0.303 0.425 0.106 0.220
BLEEP [xie_2024] 23’ 0.367 0.470 0.106 0.221 0.289 0.406 0.104 0.260 0.298 0.415 0.114 0.229
TRIPLEX [chung_2024] 24’ 0.351 0.464 0.136 0.241 0.282 0.407 0.155 0.356 0.285 0.410 0.118 0.282
@@
In comparison with existing works, as shown in Tab. 1, Gene-DML consistently achieves the best performance across all metrics, MSE, PCC(A), and PCC(H), demonstrating its superiority in accurately capturing spatial gene expression patterns. Specifically, for HER2ST dataset, Gene-DML obtains the lowest MSE of $0.210$, outperforming the second-best TRIPLEX ($0.228$) by $0.018$. In terms of PCC(A), our method achieves $0.331$, which surpasses TRIPLEX ($0.314$) by +0.017+0.017, and shows the most significant gain in PCC(H) at $0.541$, improving by +0.044+0.044 over TRIPLEX ($0.497$). As for STNet dataset, Gene-DML achieves the best MSE of $0.179$, a $0.002$ reduction from Hist2ST ($0.181$), and attains a notable gain in PCC(A) with $0.237$, improving upon TRIPLEX ($0.206$) by +0.031+0.031. In high-predictive gene correlation (PCC(H)), our method leads $0.384$, ahead $0.353$ of M2OST by +0.031+0.031. On the skin dataset skinST, Gene-DML achieves the most substantial improvements. Its MSE of $0.237$ is better than $0.268$ of TRIPLEX by +0.031+0.031. In PCC(A), it scores $0.433$, significantly outperforming the second-best TRIPLEX ($0.374$) by +0.059+0.059. For PCC(H), Gene-DML leads with $0.520$, an improvement of +0.030+0.030 over M2OST and TRIPLEX ($0.490$). Due to space limitations, all MAE results are provided in the supplementary material.
@@ 4.3 Generalization Performance of Gene-DML
Tab. 2 presents the generalization performance of Gene-DML compared to recent existing works across $3$ unseen Visium ST datasets. The results are produced by using the pre-trained Gene-DML model on the breast cancer dataset (HER2ST). Our method achieves the highest Pearson Correlation Coefficient (PCC) on both PCC(A) and PCC(H) genes across all datasets, highlighting its superior ability to capture biologically meaningful patterns that generalize beyond training distributions. Notably, Gene-DML achieves a new state-of-the-art PCC(A) of $0.160$, $0.192$, $0.155$ and PCC(H) of $0.415$, $0.411$, $0.393$ on breast Visium-1, -2, and -3, respectively, surpassing all prior approaches. While our MSE and MAE scores are not always the best, this can be attributed to Gene-DML’s emphasis on semantic alignment and structure-preserving embeddings rather than purely minimizing per-spot regression loss. The slight trade-off in MSE/MAE is acceptable, as PCC offers a more interpretable measure of functional correlation, especially for HPGs critical in downstream biological interpretation. The result of the additional $2$ skin Visium data (NCBI463 and NCBI464) is included in the supplementary.
@@ 4.4 Visualization of Cancer Biomarker Genes
GNAS [cancers14225480] and FASN [jiang2021met] are two of the highly predictive genes across all datasets we utilized, and well-established cancer biomarker genes. Fig. 3 presents a qualitative comparison of gene expression prediction with indicating PCC values for GNAS and FASN across two representative tissue WSIs from various patients, with the original WSI, gene expression ground truth, TRIPLEX prediction, and Gene-DML prediction. Specifically, each hexagonal tile corresponds to a spatial transcriptomics (ST) spot, with its color indicating the predicted gene expression level, normalized to the range $[0,1]$. Gene-DML consistently outperforms the model TRIPLEX with the second-best quantitative results in Tab. 1 in accurately recovering the spatial expression patterns, as evidenced by the visibly sharper regional expression and higher PCC values (e.g., $0.608$ vs. $0.440$ and $0.846$ vs. $0.813$ for GNAS; $0.772$ vs. $0.621$ and $0.810$ vs. $0.760$ for FASN) for both relatively hard (sample #1) and easy (sample #2) cases. The robust prediction of GNAS and FASN not only reflects the strength of features learned by the dual multi-scale discrimination but also is widely implicated in tumorigenesis and cancer progression, further underscoring the biological relevance and clinical utility of the Gene-DML framework for spatial biomarker discovery. Visualizations of additional samples are provided in the supplementary material.
Multi-scale
I-L Dis.
Global
Rep.
Cross-level
I-G Dis.
Feature
Grouping
@@
Table 3: Ablation study on framework components on the dataset skinST. Abbreviations in this table include instance-level discrimination (I-L Dis.), global-scale representation (Global Rep.), and instance-group discrimination (I-G Dis.).
@@ 4.5 Ablation Study
We performed extensive experiments to assess the effectiveness of each framework module with its key component, values of hyperparameters $k$ (feature group size), $\lambda$ (the weight of instance-group alignment loss), and $\tau_{I-G}$ (the temperature of instance-group similarity calculation).
We report the ablation results on each key component of Gene-DML, as illustrated in Tab. 3. This underscores the effectiveness of the multi-scale instance-level discrimination module and the cross-level instance-group discrimination module, together with their corresponding key components: the global-scale representation and feature grouping. Incorporating the multi-scale instance-level discrimination improves performance $0.015$ on MSE, +0.039+0.039 on PCC(A), and +0.053+0.053 on PCC(H). Similarly, adding cross-level instance-group discrimination yields a gain $0.017$ on MSE, $0.051$ on PCC(A), and $0.082$ on PCC(H) from the plain bimodal framework. When both modules are integrated, Gene-DML achieves the best performance across all metrics, showing a cumulative improvement of $0.060$ MSE, +0.164+0.164 PCC(A), and +0.124+0.124 PCC(H) over the baseline. Besides, comparative experiments with and without the global-scale representation and feature grouping further highlight their essential contribution to the effectiveness of the corresponding modules. These results validate the complementary benefits of a dual-pathway multi-level discrimination strategy.
@@
Tab. 4 demonstrates the impact of varying the number of groups $k$ in the feature grouping (Sec. 3.2.1). The model performance is relatively robust across different values but achieves the best results when $k=25$, with the lowest MSE and MAE, and the highest PCC(A) and PCC(H). These results suggest that an appropriately chosen group size provides a good balance between semantic expressiveness and discrimination, facilitating more effective instance-group alignment. Tab. 5 presents an ablation study on the impact of the weighting coefficient $\lambda$ for the instance-group contrastive loss. The results show that incorporating instance-group discrimination consistently improves performance across all metrics. Specifically, when $\lambda=0.8$, the best overall results are obtained, achieving the lowest MSE and MAE, along with the highest PCC scores. This indicates that moderately weighting the instance-group loss effectively enhances cross-modal instance-group alignment without overwhelming the training objective. Tab. 6 shows the effect of the temperature parameter $\tau_{I\text{-}G}$ used in the instance-group similarity calculation. $\tau_{I\text{-}G}$ controls the sharpness of the softmax distribution over similarity scores. The results show that setting $\tau_{I\text{-}G}=0.07$ achieves optimal performance across all evaluation metrics. The ablation of HER2ST and STNet is included in the supplementary.
@@ 5 Conclusion
Appendix B: Additional Generalization Performance
- Provides the generalization performance of Gene-DML compared to TRIPLEX across $2$ additional unseen skin cancer Visium ST data.
Appendix C: Additional Ablation Study
@@
Appendix D: Additional Visualization
@@
Appendix E: Additional Implementation Details
@@ Appendix A Additional Cross-validation Performance
Appendix B Additional Generalization Performance
Model NCBI463 NCBI464
[…]
TRIPLEX 0.342 0.478 0.103 0.360 0.416 0.538 0.091 0.261
Gene-DML (ours) 0.253 0.413 0.109 0.377 0.292 0.445 0.140 0.441
Table 8: Generalization performance on NCBI463 and NCBI464.
We further extend the evaluation by adding two skin cancer Visium datasets (NCBI463 and NCBI464) [Jaume_2025] to verify the skinST pre-trained models. Tab. 8 further underscores Gene-DML’s outstanding generalization performance, achieving state-of-the-art results across all evaluation metrics MSE, MAE, PCC(A), and PCC(H).
Appendix C Additional Ablation Study
To evaluate the dual-pathway multi-level discrimination framework, Gene-DML, we perform an ablation study by adding each discrimination pathway sequentially together with their corresponding key components: the global-scale representation and feature grouping, and report the results of MSE, MAE, PCC(A), and PCC(H) metrics. Tab. 9 and Tab. 13 demonstrate the effectiveness of each component in Gene-DML on datasets HER2ST and STNet. They show that, by removing both pathways, a plain bimodal framework yields the weakest performance. By incorporating the multi-scale instance-level discrimination pathway and cross-level instance-group discrimination pathway, respectively, a significant improvement is achieved.
@@ Appendix C Additional Ablation Study
Table 9: Ablation study on components of the dataset HER2ST. Abbreviations in this table include instance-level discrimination (I-L Dis.), global-scale representation (Global Rep.) and instance-group discrimination (I-G Dis.).
@@
Table 13: Ablation study on components of the dataset STNet. Abbreviations in this table include instance-level discrimination (I-L Dis.), global-scale representation (Global Rep.) and instance-group discrimination (I-G Dis.).
@@
Appendix D Additional Visualizations
In this section, we provide additional visualizations of the gene expression prediction results, which are normalized into the range $[0,1]$, of the breast cancer biomarker genes GNAS and FASN. Fig. 4 demonstrates the spatial gene expression prediction for GNAS and FASN on the HER2ST dataset. Each row displays a WSI sample alongside the ground-truth expression map and predictions from TRIPLEX (the method achieved the second-best result evaluated in Tab. 1 of Sec. 4.2.) and the proposed Gene-DML. Across both genes and multiple WSI samples, Gene-DML consistently produces spatial expression patterns that are visually closer to the ground truth and quantitatively higher in performance, as reflected by the reported Pearson correlation coefficients in parentheses. The same conclusion holds for Fig. 5, which shows the spatial gene expression prediction results on the dataset STNet, further validating the strong gene expression prediction capacity of Gene-DML.
@@ Appendix D Additional Visualizations
Appendix E Additional Implementation Details
To ensure fairness, following TRIPLEX [chung_2024], we utilize the various $250$ selected genes for each dataset, i.e., HER2ST, STNet, and skinST, that are related to morphology in histopathology to evaluate the performance of gene expression prediction. For the results of the MSE, MAE, and PCC(A) metrics in the cross-validation experiments, the same $250$ genes were utilized for both the training and testing processes. Fig. 6, Fig. 7, and Fig. 8 demonstrate the $250$ genes exploited to predict gene expression profiles on datasets HER2ST, STNet, and skinST, respectively.